OpenStudy (anonymous):

78= -2(m+3) +m can someone help me out?

sam (.sam.):

Since m is on the right-hand side of the equation, switch the sides so it is on the lefthand side of the equation -2(m+3)+m=78 multiply -2 by each term inside the parentheses -2m-6+m=78 Since -2m and m are like terms, subtract m from -2m -m-6=78 Move all terms not containing m to the right side of the equation. -m=84 multiply -m by -1 m=84*-1 Divide each term in the equation by -1. m=-84

OpenStudy (anonymous):

78 = -2(m+3) + m 78 = m(-2) + 3(-2) + m 78 = -2m - 6 + m 78 + 6 = -2m + m 84 = -m m = -84

OpenStudy (anonymous):

78=-2(m-3)+m I think it looks neater when (m+84)=0, where m=-84